OpenStudy (anonymous):

Evaluate the expression by using the given values of the variables. y ÷ 4 • 3 – 6 ÷ x + 8 where x = 3 and y = 24

OpenStudy (anonymous):

Just plug x in where it should go and the same for y and solve

OpenStudy (anonymous):

@iGreen @joejoe213

OpenStudy (anonymous):

@jabez177

OpenStudy (anonymous):

\[24\div4\times6\div3+8\]

OpenStudy (igreen):

\(y \div 4 \times 3 - 6 \div x + 8\) Plug in y and x: \(24 \div 4 \times 3 - 6 \div 3 + 8\) Divide: \(6 \times 3 - 6 \div 3 + 8\) Multiply: \(18 - 6 \div 3 + 8\) Divide: \(18 - 2 + 8\) Okay, can you finish that? Just subtract, then add.

OpenStudy (anonymous):

24

OpenStudy (igreen):

Yep, that's your answer!